# Latent Action Diffusion for Cross-Embodiment Manipulation

### v2

@@ Abstract
End-to-end learning approaches offer great potential for robotic manipulation, but their impact is constrained by data scarcity and heterogeneity across different embodiments. In particular, diverse action spaces across different end-effectors create barriers for cross-embodiment learning and skill transfer. We address this challenge through diffusion policies learned in a latent action space that unifies diverse end-effector actions. We first show that we can learn a semantically aligned latent action space for anthropomorphic robotic hands, a human hand, and a parallel jaw gripper using encoders trained with a contrastive loss. Second, we show that by using our proposed latent action space for co-training on manipulation data from different end-effectors, we can utilize a single policy for multi-robot control and obtain up to 25% improved manipulation success rates, indicating successful skill transfer despite a significant embodiment gap. Our approach using latent cross-embodiment policies presents a new method to unify different action spaces across embodiments, enabling efficient multi-robot control and data sharing across robot setups. This unified representation significantly reduces the need for extensive data collection for each new robot morphology, accelerates generalization across embodiments, and ultimately facilitates more scalable and efficient robotic learning.
Keywords: Imitation Learning, Cross-Embodiment Learning, Manipulation
Figure 1: We introduce a framework for learning cross-embodied manipulation policies through latent action diffusion. First, a shared latent action space is discovered through contrastive learning from pairs of aligned end-effector poses. Training diffusion policies with latent actions enables multi-embodiment control with a single policy and realizes skill transfer between embodiments.
1 Introduction
Robotic manipulation holds vast transformative potential to address global labor shortages through easy-to-deploy robotic workers able to adapt to different settings. End-to-end learning of manipulation policies is set to equip robots with the necessary skill set and intelligence. The end-to-end learning paradigm has proven its success in high-data regimes for language and vision models. Robot learning, however, presents novel challenges: imitation learning models are still in a data-bound regime where real-world performance is largely dictated by the volume and diversity of the training data. Scaling up both of these factors inevitably requires pooling together data from different robotic embodiments. However, while it is possible to increase data volume and diversity through cross-embodiment learning, the heterogeneity across observation and action spaces of different robotic embodiments poses significant barriers for skill transfer across embodiments (the “embodiment gap“).
@@ 1 Introduction
In particular, we focus on embodiment transfer among single-arm robots with different end-effectors. For our experiments, we utilize the Faive robotic hand [8], the mimic hand [9] and a Franka parallel gripper. In two experiments, pairing data from each dexterous hand with data from the Franka gripper utilizing our proposed framework (Figure 1), we show both cross-embodiment control with a single policy and unveil positive skill transfer across both embodiments with up to 13% performance (10.16% average) improvement obtained through co-training, despite a significant embodiment gap.
@@
2 Related Works
@@
Learning from cross-embodiment data is a promising path towards scaling up both the volume and diversity of training data for robot policies. Brohan et al. [1] showed positive skill transfer by co-training on multi-robot datasets with the same action space. Building on the Open-X-Embodiment collaboration [10], multiple approaches have explored large-scale pretraining on more diverse robot data [2, 3, 11]. However, these works rely on constraining the action space for pretraining to a 7-dimensional space for single-arm robots with a parallel jaw gripper, discarding pretraining data with other action spaces. Shaw et al. [4], Wang et al. [7] and Kareer et al. [6] have investigated pretraining on human video data by extracting human actions from video and retargeting them to the representation required by the action space of their respective robots. Doshi et al. [12] use multiple action heads, each representing a distinct action space, to pretrain on more diverse data mixtures. Black et al. [13] pretrain on different action spaces by aligning shared features in the action space and padding unused feature dimensions. Ye et al. [14] and Chen et al. [15] use representations of video frame prediction models as coarse latent action representations to pretrain policies on latent action spaces learned from unlabelled video data.
Previous works relied either on limited alignment strategies (e.g., retargeting human actions to robot actions) that are not scalable across multiple embodiments, or used end-to-end learning approaches without explicit action space alignment. In contrast, our approach provides a framework that can scale to multiple end-effectors while providing a unified action representation that can provide useful inductive biases for manipulation policies, which is not guaranteed with end-to-end learning approaches.
Figure 2: Overview of our proposed framework for semantic alignment of end-effector actions via retargeting (stage 1), training contrastive encoders and decoders for the learned latent action space (stages 2 and 3) and policy learning in latent space (stage 4). We validate our methodology for latent spaces learned with data from the Faive hand, mimic hand, human hands and a Franka gripper and data from the mimic hand and a Franka gripper.
3 Methodology
@@
We propose a four-stage framework for cross-embodied latent space imitation learning (Figure 2). The objective of our framework is to learn policies for single-arm robots with different end-effectors in a unified latent action space. The key insight is that learning aligned representations for different end-effector action spaces can be viewed as a multimodal representation learning problem. Based on this perspective, we design a pipeline for cross-embodied latent imitation learning comprised of the following steps: generating paired action data, learning encoders and decoders for the shared latent space, and latent policy learning.
3.1 Creating Aligned Action Pairs
@@
Multimodal representation learning architectures for $M$ modalities generally rely on tuples containing paired data of the form $\mathbf{x}_{i}=\left(x_{i}^{1},x_{i}^{2},\dots,x_{i}^{M}\right)$, where there is some form of cross-modal correspondence between the elements of each tuple. In multimodal learning, correspondences between data modalities are typically created through manual annotation (e.g., image-caption pairs [16]) or created with modality-specific expert models (e.g., creating depth pseudolabels from RGB images [17]). In the context of robotics, we are looking for alignment functions between different action spaces which allow us to establish mappings in between the action spaces. We focus on aligning action spaces of different end-effectors (human hands, anthropomorphic robotic hands, parallel jaw gripper, …). For this subproblem, retargeting functions from human hands to robotic end-effectors are a useful prior for alignment, as they typically already exist for different embodiments in order to teleoperate robots.
@@
3.1.1 Action Representations
For human hands, we derive a 189-dimensional pose representation $\mathbf{\theta}_{H}$ using the local transformations in between the 21 joints according to the kinematic chain of the hand. To represent rotations, we utilize the continuous 6D rotation representation proposed by Zhou et al. [18]. Poses for the Faive hand or mimic hand are represented as an 11- or 16-dimensional vector of joint angles $\mathbf{\theta}_{F}$ or $\mathbf{\theta}_{M}$ respectively. Poses of parallel jaw grippers are represented as normalized one-dimensional gripper width $\theta_{P}\in[0,1]$.
3.1.2 Retargeting
For retargeting, we follow the technique introduced by Sivakumar et al. [19], which utilizes keyvectors for both the human and robot hand. The keyvectors $v_{i}^{\{H,M\}}(\mathbf{\theta}_{\{H,M\}})$ are vectors from the palm to each fingertip and from each fingertip to all other fingertips and provide a unifying representation that can be defined for any hand with a notion of fingertips. To map from human hands to complex robotic hands such as the mimic hand, we can formulate retargeting functions as a minimum-energy solution to the squared keyvector difference of the human and the robot hand. By using the forward kinematics of each hand, we can determine the keyvectors as a function of its respective pose representation $\mathbf{\theta}_{H}$ or $\mathbf{\theta}_{H}$. As a concrete example, to retarget from a human hand pose $\mathbf{\theta}_{H}$ to a mimic hand pose $\mathbf{\theta}_{M}$, we can directly optimize over the $\mathbf{\theta}_{M}$ with the differentiable objective shown in Equation 2. Each pair of keyvectors has a scaling factor $s_{i}$, which is used to compensate for different finger lengths. For all 15 keyvectors, scaling factors are determined through qualitative evaluation. The resulting retargeting function can be expressed as follows:
$\mathbf{\theta}_{M}(\mathbf{\theta}_{H})=\text{argmin}_{\mathbf{\theta}_{F}}\sum_{i=1}^{15}\left|\left|v_{i}^{H}(\mathbf{\theta}_{H})-s_{i}v_{i}^{F}(\mathbf{\theta}_{M})\right|\right|_{2}^{2}$ (2)
@@ 3.1.2 Retargeting
$\theta_{P}(\mathbf{\theta}_{H})=\min_{\theta_{P}}\left(\min_{i}\frac{\left|\left|v_{i}^{H}(\mathbf{\theta}_{H})\right|\right|}{W},1\right)$ (3)
@@
3.2 Contrastive Latent Space Learning
For a shared latent action space, it is crucial that 1) for each modality, sufficient information is encoded such that we can precisely reconstruct end-effector poses and 2) the latent space has a coherent structure, meaning that the cross-modal alignment present in the model inputs during training is upheld in the learned latent space. To achieve both of these goals, we propose a two-step learning procedure: first, using batches with $B$ aligned end-effector poses, $M$ modality-specific encoders $q_{m},m\in 1\dots M$ are trained that project actions $x_{m}$ from each input modality into a shared latent space, where we utilize a pairwise InfoNCE loss [20] to ensure alignment within the batch:
$\mathcal{L}_{\text{contrastive}}=\frac{1}{M(M-1)}\sum_{i=1}^{M}\sum_{j=i+1}^{M}\left(-\frac{1}{B}\sum_{n=1}^{B}\log\frac{\exp(q_{i}(x_{i}^{n})\cdot q_{j}(x_{j}^{n})/\tau)}{\sum_{k=1}^{B}\exp(q_{i}(x_{i}^{n})\cdot q_{j}(x_{j}^{k})/\tau)}\right)$ (4)
@@ 3.2 Contrastive Latent Space Learning
$\mathcal{L}_{\text{recon}}$ $=\frac{1}{M}\sum_{i=1}^{M}\sum_{n=1}^{B}\left|\left|p_{i}\left(q_{i}\left(x_{i}^{n}\right)\right)-\hat{x}_{i}^{n}\right|\right|_{2}^{2}$ (5)
@@
Hyperparameters for training can be found in Section 7.3. Qualitative visualizations are shown in Figure 3. An ablation study evaluating the impact of jointly finetuning the encoders and temperature annealing is shown in Table 1.
3.3 Policy Learning
With a learned latent action space, we employ Diffusion Policy [21] to map shared observations across datasets to latent actions of different embodiments (Fig. 4). We validate our method across two diffusion policy implementations: a transformer-based implementation [22] and a U-Net-based implementation [21]. The architecture for the transformer-based policy is visualized in detail in Figure 4. The transformer-based policy is utilized in experiments with the Faive hand and Franka gripper, whereas the U-Net-based implementation is used for experiments with the mimic hand and the Franka gripper. Both architectures are described in more detail in the appendix in Section 7.2.
Figure 4: Detailed overview over the policy network architecture for an action chunk size $k$, action dimension $d_{\text{act}}$, an observation history $T_{\text{hist}}$, a token dimension $d_{\text{token}}$, a batch size $B$, and $N_{\text{enc}}$ transformer encoder layers. The observations are fixed as global conditioning during the diffusion process whereas the noisy latent actions are updated with each denoising step.
Table 1: Ablation study of the contrastive action model components.
Model Configuration SR-Loss CR-Loss
@@ 3.3 Policy Learning
Full Model (ours) 0.762 3.7e-8 0.002 214.20
w/o Temperature Annealing (TA) 0.948 1.5e-8 0.007 286.64
w/o Finetuning (FT) 44.76 2.6e-8 0.013 391.85
w/o FT and TA 49.765 2.1e-8 0.02 397.23
4 Experimental Results and Discussion

### v3

# ]references.bib

@@ Abstract
End-to-end learning is emerging as a powerful paradigm for robotic manipulation, but its effectiveness is limited by data scarcity and the heterogeneity of action spaces across robot embodiments. In particular, diverse action spaces across different end-effectors create barriers for cross-embodiment learning and skill transfer. We address this challenge through diffusion policies learned in a latent action space that unifies diverse end-effector actions. We first show that we can learn a semantically aligned latent action space for anthropomorphic robotic hands, a human hand, and a parallel jaw gripper using encoders trained with a contrastive loss. Second, we show that by using our proposed latent action space for co-training on manipulation data from different end-effectors, we can utilize a single policy for multi-robot control and obtain up to 25.3% improved manipulation success rates, indicating successful skill transfer despite a significant embodiment gap. Our approach using latent cross-embodiment policies presents a new method to unify different action spaces across embodiments, enabling efficient multi-robot control and data sharing across robot setups. This unified representation significantly reduces the need for extensive data collection for each new robot morphology, accelerates generalization across embodiments, and ultimately facilitates more scalable and efficient robotic learning.
I Introduction
End-to-end learning is a promising path toward creating adaptable, generalist robots. Scaling up both the data volume and diversity to match the desired model capabilities is extremely resource-intensive and expensive, and inevitably requires pooling together data from different robotic embodiments. However, efficiently learning from different embodiments remains a significant challenge, as observation and action spaces vary significantly across robots (the “embodiment gap“).
@@ I Introduction
In particular, we focus on embodiment transfer among single-arm robots with different end-effectors. For our experiments, we utilize the Faive robotic hand [8], the mimic hand [9] and a Franka parallel gripper. In two experiments, pairing data from each dexterous hand with data from the Franka gripper utilizing our proposed framework (LABEL:fig:title_figure), we compare latent diffusion policies co-trained on cross-embodiment data with single-embodiment diffusion policies. We demonstrate that our methodology enables both cross-embodiment control with a single policy and facilitates positive skill transfer, with up to 25.3% performance (13.4% average) improvement compared to single-embodiment diffusion policies, despite a significant embodiment gap.
@@
Figure 1: The three-stage process for learning the cross-embodiment latent action space. Stage 1: Aligned end-effector (EEF) poses are generated by retargeting human hand poses to different robot end-effectors. Stage 2: Embodiment-specific encoders are trained to project these actions into a shared latent space using a contrastive loss. Stage 3: Decoders are trained to reconstruct the original poses from the latent space, and encoders are fine-tuned to improve reconstruction quality.
I-A Contributions
Our main contributions are:
  1. 1.
    We introduce a general framework that unifies diverse end-effector action spaces into a single, semantically aligned latent space using contrastive learning, enabling downstream learning across diverse robotic embodiments.
  2. 2.
    We show that factorizing diffusion policies into a latent, embodiment-agnostic policy and embodiment-specific action decoders enables multi-robot control across substantially different robot morphologies.
  3. 3.
    We demonstrate substantial real-world performance gains — up to 25.3% success rate improvement — from cross-embodiment co-training, and provide ablation studies validating our architecture for learning latent action spaces.
II Related Works
Learning from cross-embodiment data is a promising path towards scaling up both the volume and diversity of training data for robot learning that has been considered from various angles in prior works.
Constrained Action Space
@@ II Related Works
[1] showed positive skill transfer by co-training on multi-robot datasets with the same action space. Building on the Open-X-Embodiment collaboration [10], multiple approaches have explored large-scale pretraining on more diverse robot data [2, 3, 11]. However, these works rely on constraining the action space for pretraining to a 7-dimensional space for single-arm robots with a parallel jaw gripper, discarding any pretraining data with other action spaces such as dexterous hands.
One-Way Retargeting
[4], [7] and [6] have investigated pretraining on human video data by extracting human actions from video and retargeting them to the explicit action representation required by their robots. As opposed to using explicit actions, our methodology uses a learned action space. Using a learned action space instead of explicit actions provides more expressive action representations for diverse robotic embodiments, and enables multi-robot control.
Alignment-Free Methods
Multiple works have investigated cross-embodiment learning without explicit alignment between different action spaces, using large-scale pretraining to obtain useful representations for robotic manipulation. [12] assign different action spaces to separate feature dimensions of the model output and padding unused action spaces for each dataset. [13] use multiple action heads, each representing a distinct action space, to pretrain on diverse data mixtures. [14] and [15] use representations of video frame prediction models as coarse latent action representations to pretrain policies on latent action spaces learned from unlabelled video data.
However, both sample-efficiency and skill transfer from and to the high-dimensional, complex action spaces of dexterous robotic end-effectors remain open questions with alignment-free approaches. Our method, utilizing a unified latent action space, enables skill transfer between highly diverse embodiments with comparatively few samples.
Each of these prior approaches makes different trade-offs between action space expressiveness, alignment requirements, and data efficiency. In contrast to prior work, our framework leverages retargeting not as a direct mapping to a fixed action represetation for policy learning, but rather as a strong inductive bias to structure a unified latent action space for policy learning. This approach avoids the limitations of constrained action spaces and rigid one-way mappings, allowing it to unify highly diverse end-effectors, including dexterous hands. By building this semantic alignment into the action representation used for policy learning, our method facilitates sample-efficient skill transfer across significant embodiment gaps, addressing a key challenge faced by large-scale, alignment-free approaches.
III Methodology
Our proposed framework for cross-embodied policy learning consists of two parts: learning a latent action space (Figure 1) and training latent policies. In this work, we focus on policy learning for single-arm robots with different end-effectors in a unified latent action space.
@@ III Methodology
Our key insight is that learning aligned representations for different end-effector action spaces can be viewed as a multimodal representation learning problem. Based on this perspective, we design a pipeline for cross-embodied latent imitation learning comprised of the following steps: generating paired action data, learning encoders and decoders for the shared latent space, and latent policy learning.
III-A Creating Aligned Action Pairs
@@
Multimodal representation learning architectures for $M$ modalities generally rely on tuples containing paired data of the form $\mathbf{x}_{i}=\left(x_{i}^{1},x_{i}^{2},\dots,x_{i}^{M}\right)$, where there is some form of cross-modal correspondence between the elements of each tuple. In multimodal learning, correspondences between data modalities are typically created through manual annotation (e.g., image-caption pairs [16]) or created with modality-specific expert models (e.g., creating depth pseudolabels from RGB images [17]).
In the context of robotics, we are looking for alignment functions between different action spaces which allow us to establish mappings in between the action spaces. We focus on aligning action spaces of different end-effectors (human hands, anthropomorphic robotic hands, parallel jaw gripper, …). For this subproblem, retargeting functions from human hands to robotic end-effectors are a useful prior for alignment, as they typically already exist for different embodiments in order to teleoperate robots.
@@
III-A1 Action Representations
For human hands, we derive a 189-dimensional pose representation $\mathbf{\theta}_{H}$ using the local transformations in between the 21 joints according to the kinematic chain of the hand. To represent rotations, we utilize the continuous 6D rotation representation proposed by [18]. Poses for the Faive hand or mimic hand are represented as an 11- or 16-dimensional vector of joint angles $\mathbf{\theta}_{F}$ or $\mathbf{\theta}_{M}$ respectively. Poses of parallel jaw grippers are represented as normalized one-dimensional gripper width $\theta_{P}\in[0,1]$.
III-A2 Retargeting
For retargeting, we follow the technique introduced by [19], which utilizes keyvectors for both the human and robot hand. The keyvectors $v_{i}^{\{H,M\}}(\mathbf{\theta}_{\{H,M\}})$ are vectors from the palm to each fingertip and from each fingertip to all other fingertips and provide a unifying representation that can be defined for any hand with a notion of fingertips. To map from human hands to complex robotic hands such as the mimic hand, we can formulate retargeting functions as a minimum-energy solution to the squared keyvector difference of the human and the robot hand. By using the forward kinematics of each hand, we can determine the keyvectors as a function of its respective pose representation $\mathbf{\theta}_{H}$ or $\mathbf{\theta}_{H}$. As a concrete example, to retarget from a human hand pose $\mathbf{\theta}_{H}$ to a mimic hand pose $\mathbf{\theta}_{M}$, we can directly optimize over the $\mathbf{\theta}_{M}$ with the differentiable objective shown in Equation 2. Each pair of keyvectors has a scaling factor $s_{i}$, which is used to compensate for different finger lengths. For all 15 keyvectors, scaling factors are determined through qualitative evaluation. The resulting retargeting function can be expressed as follows:
[…]
@@ III-A2 Retargeting
[…]
@@
III-B Contrastive Latent Space Learning
For a shared latent action space, it is crucial that 1) for each modality, sufficient information is encoded such that we can precisely reconstruct end-effector poses and 2) the latent space has a coherent structure, meaning that the cross-modal alignment present in the model inputs during training is upheld in the learned latent space. To achieve both of these goals, we propose a two-step learning procedure: first, using batches with $B$ aligned end-effector poses that were generated via retargeting, $M$ modality-specific encoders $q_{m},m\in 1\dots M$ are trained that project actions $x_{m}$ from each input modality into a shared latent space, where we utilize a pairwise InfoNCE loss [20] to ensure alignment within the batch:
$\begin{split}&\mathcal{L}_{\text{contrastive}}=\frac{1}{M(M-1)}\\ &\sum_{i=1}^{M}\sum_{j=i+1}^{M}\left(-\frac{1}{B}\sum_{n=1}^{B}\log\frac{\exp(q_{i}(x_{i}^{n})\cdot q_{j}(x_{j}^{n})/\tau)}{\sum_{k=1}^{B}\exp(q_{i}(x_{i}^{n})\cdot q_{j}(x_{j}^{k})/\tau)}\right)\end{split}$ (4)
@@ III-B Contrastive Latent Space Learning
[…]
@@
We use a cross-reconstruction (CR) loss for validation. From modality $i$ to $j$, given paired end-effector poses $(x_{i}^{n},x_{j}^{n})$, the CR-Loss is:
@@
We encode data from modality $x_{i}^{n}$, decode it to modality $j$ and evaluate the result versus the paired ground truth data $x_{j}^{n}$. The self-reconstruction (SR) loss used for validation is $\mathcal{L}_{\text{SR(i)}}=\mathcal{L}_{\text{CR(i,i)}}$.
Action decoders and encoders are parameterized by standard multi-layer-perceptrons (MLPs) as encoders and decoders (Figure 1). The input layer consists of a linear layer, ReLU layer and a dropout layer. For each hidden layer, we first have a normalization layer, then a linear layer, followed by a ReLU (rectified linear unit) activation and a dropout layer. After the hidden layers, a last linear layer is used to project the output to the desired dimension.
III-C Policy Learning
With a learned latent action space, we employ Diffusion Policy [21] to map shared observations across datasets to latent end-effector actions of different embodiments and non-latent wrist poses of the robot arm as shown in LABEL:fig:title_figure. The encoders and decoders of the contrastive action model stay frozen for latent policy learning, with the diffusion objective being applied on the denoised latent end-effector actions and wrist poses. For inference, the suitable embodiment-specific decoder is used to obtain a decoded end-effector pose from the output of the latent policy.
We validate our method across two diffusion policy implementations: a transformer-based implementation [22] and a U-Net-based implementation [21]. The transformer-based policy is utilized in experiments with the Faive hand and Franka gripper, whereas the U-Net-based implementation is used for experiments with the mimic hand and the Franka gripper. Specifications for the models used in the experiments are given in Table I.
For co-training on differently sized datasets, we assign normalized weights $w_{j}$ to all datasets. During training, we seek to combine samples from all datasets to fill batches with $B$ samples in total. We sample per-dataset sub-batches with appropriately rounded sizes $\text{round}(\frac{B}{w_{j}})$, project the actions into the shared latent action space, normalize the sub-batches, and then concatenate them into a single batch for efficient training. Through this mechanism, the weight of each dataset approximately represents a sampling probability for each training step. In our experiments, all datasets have the same weights, such that the model is exposed to equal amounts of data from all embodiments.
TABLE I: Diffusion Policy Specifications
Hyperparameter Transformer U-Net
Parameter Count $\sim$40M $\sim$44M
Vision Encoder Resnet18 ViT-S
Batch Size 300 256
Horizon 21 timesteps 48 timesteps
Diffusion Noise Schedule Squared cosine Squared cosine
$\beta_{\text{start}}$ 0.0001 0.0001
$\beta_{\text{end}}$ 0.02 0.02
Peak Learning Rate 0.0001 0.0001
Optimizer AdamW AdamW
Training Steps 90k 40k
TABLE II: Contrastive Action Model Specifications
Hyperparameter Value
Parameter Count (single encoder/decoder) $\sim$26k
Batch Size 16384
Learning Rate 0.00001
Finetuning Learning Rate 0.0001
Optimizer AdamW
Weight Decay 0.001
Temperature Schedule Exponential
[…]
Latent Space Dimension 16
$\lambda$ 0.1
MLP Hidden Dimensions 32, 128, 128, 32
Training Epochs (Encoders) 5000
Training Epochs (Decoders + finetune encoders) 10000
TABLE III: Ablation study on the impact of temperature annealing (TA) and finetuning (FT) on the contrastive action model, evaluating self- and cross-reconstruction losses.
Method SR-Loss CR-Loss
$mimic\rightarrowFranka Franka\rightarrowmimic$
Full (ours) 0.762 3.7e-8 0.002 214.20
no TA 0.948 1.5e-8 0.007 286.64
no FT 44.76 2.6e-8 0.013 391.85
no FT&TA 49.765 2.1e-8 0.02 397.23
Figure 3: Success rates for three different tasks comparing single-embodiment diffusion policies to cross-embodied latent diffusion policies trained on data from both embodiments for each task. Block stacking: 200 demos per embodiment, one external camera. Block pick and place: 200 demos per embodiment, one external camera + wrist camera for mimic hand, replaced by zero-padding for Franka gripper. Plush toy pick and place: 100 demos per embodiment, one external camera.
Figure 4: Cross-embodiment policy rollouts for two pick and place tasks in different settings. The robots in each respective setting (left: mimic hand, middle: mimic hand, Franka gripper, right: Faive hand, Franka gripper) are controlled by a single cross-embodiment diffusion policy, demonstrating multi-robot control.
IV Experimental Results and Discussion